OpenStudy (anonymous):

can someone help me solve this 3 part word problem. 2. As you are leaving the community, you notice another new community just down the street. Because you are in the area, you decide to inquire about it. a. The sales representative here tells you they also have two floor plans available, but they only have 38 homes still for sale. Write an equation that illustrates the situation. Use x and y to denote floor plan one and floor plan two respectively. b. The representative tells you that floor plan one sells for $175,000 and floor plan two sells for $200,000. She also mentions that all the a

OpenStudy (anonymous):

The complete question ?

OpenStudy (anonymous):

that is the complete question

OpenStudy (anonymous):

No there is one figure missing at the end

OpenStudy (anonymous):

It is x + y = 36 and 175000 x + 200000y = ?

OpenStudy (anonymous):

for which one a, b, c,

OpenStudy (anonymous):

The question Mark is the missing figure

OpenStudy (anonymous):

In b

OpenStudy (anonymous):

can you help me with C

OpenStudy (anonymous):

There is no c in the question I can see

OpenStudy (anonymous):

c. Use elimination to determine how many houses are available in each floor plan. Explain how you arrived at your answer.

OpenStudy (anonymous):

I'm missing the last figure in b. Can you re type b please

OpenStudy (anonymous):

b. The representative tells you that floor plan one sells for $175,000 and floor plan two sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in Part a.

OpenStudy (anonymous):

i was missing the 7200000 in the question Hang on....

OpenStudy (anonymous):

2. As you are leaving the community, you notice another new community just down the street. Because you are in the area, you decide to inquire about it. a. The sales representative here tells you they also have two floor plans available, but they only have 38 homes still for sale. Write an equation that illustrates the situation. Use x and y to denote floor plan one and floor plan two respectively. b. The representative tells you that floor plan one sells for $175,000 and floor plan two sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in Part a. c. Use elimination to determine how many houses are available in each floor plan. Explain how you arrived at your answer.

OpenStudy (anonymous):

X + y = 36 175 x + 200y = 7200 X = y - 36 175y - 6300 + 200y = 7200 375y = 13500 y = 36 they are all floor plan 2

OpenStudy (anonymous):

thanks is this the complete answer